OpenStudy (anonymous):

the ration of two numbers is 1 to 3, and the sum of the numbers is 28. What is the smaller number?

OpenStudy (radar):

This is a problem where you just write down what is given and solve. It is given that a ratio of 1 to 3 or y/x =3 or y = 3x Now the sum of the two numbers is given x + y = 28. Now just substitute the 3x for the y getting: x + 3x = 28 or 4x = 28. Now solve for x then solve for y.

OpenStudy (anonymous):

So x is 7 then whats Y? thank you by the way.

OpenStudy (radar):

Good, just continue to solve, you were told that x + y = 28. You have successfully solve for x giving it a value of 7. Now substitute 7 for x in the x + y = 28 and solve for y.

OpenStudy (radar):

Algebra is power. Good luck with it.